OpenStudy (anonymous):

Tom has a collection of 30 CDs and Nita has a collection of 18 CDs. Tom is adding 1 CD a month to his collection while Nita is adding 5 CDs a month to her collection. Find the number of months after which they will have the same number of CDs.

OpenStudy (anonymous):

Write an equation for this: 30 + x = 18 + 5x. Solve for x to get x = 3. There will be 3 months.